OpenStudy (anonymous):

my final question for the night, prove that 2cos(x+pi/4)cos(x-pi/4)=cos2x

OpenStudy (anonymous):

expand cos(x+pi/4) and cos(x-pi/4) using the identity \[\cos(a+b)=\cos a \cos b-\sin a\sin b\]

OpenStudy (anonymous):

i got that far i wrote cos2x =cosx.cos45deg-sinx.sin45deg? what next

OpenStudy (anonymous):

if 45deg does that mean all 45deg

OpenStudy (anonymous):

please help me finish this question

OpenStudy (anonymous):

sorry i lost connection u mean cos(x+pi/4) =cosx.cos45deg-sinx.sin45deg ?

OpenStudy (anonymous):

its double angle formula 2cosx+pi/4

OpenStudy (anonymous):

make that just 2cs=

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

what do u mean ? by 2cs=

OpenStudy (anonymous):

2cosx

OpenStudy (anonymous):

well let me tell u something : cos(x+pi/4) =cosx.cos45deg-sinx.sin45deg right?

OpenStudy (anonymous):

but it has to cancel from here to get cos2x?

OpenStudy (anonymous):

this equation leaves 0

hartnn (hartnn):

how do u get 0? cos 45=sin 45 = 1/(sqrt 2) this makes cos(x+pi/4) = (1/(sqrt 2))*(cos x -sin x) similarly do it for cos(x-pi/4) ....

OpenStudy (anonymous):

i don,'t know how i got it, if you know how to come up with the answer it would be good if you could show it

OpenStudy (anonymous):

i wanna do this completely...then u tell us which part is confusing.. \[\sin \frac{\pi}{4}=\cos \frac{\pi}{4}=\frac{\sqrt{2}}{2}\]so using the first identity i gave u\[\cos(x+\frac{\pi}{4})=\cos x \cos \frac{\pi}{4}-\sin x \sin \frac{\pi}{4}=\frac{\sqrt{2}}{2}(\cos x-\sin x)\]\[\cos(x-\frac{\pi}{4})=\cos x \cos \frac{\pi}{4}+\sin x \sin \frac{\pi}{4}=\frac{\sqrt{2}}{2}(\cos x+\sin x)\] so u will come up with\[2\cos(x-\frac{\pi}{4})\cos(x+\frac{\pi}{4})=2\frac{\sqrt{2}}{2}\frac{\sqrt{2}}{2}(\cos x-\sin x)(\cos x+\sin x)\\ =(\cos x-\sin x)(\cos x+\sin x)=\cos^2 x-\sin^2 x=\cos 2x\]using the fact that\[(a-b)(a+b)=a^2-b^2\]